### v1

@@ VIII Diagnosis Effectiveness
Figure 6: Test on Easy Level
@@
Figure 7: Test on Medium Level
@@
Figure 8: Test on Difficult Level
Table VI: Navigation performance under different high-resistance area up-sampling ratios $\lambda$.
@@
Table VII: Navigation performance under different trajectory vector change thresholds $\eta$.
@@
(a) Average Time
[…]
(b) Average Reward
[…]
(c) Average Success Rate
Figure 9: Average time, reward, and success rate of Digimon and APPLR for different checkpoints during training.
IX Additional Experiments
IX-A More Results of Efficiency Comparison
@@
In Figure 9, we present additional metrics for various meta-planner checkpoints during training, including average traveling time, reward, and success rate. It clearly shows that Digimon outperforms APPLR across all three metrics. Notably, Digimon requires fewer training steps to achieve shorter traveling times, highlighting the method’s efficiency. Similarly, within the same amount of training steps, Digimon can also achieve a higher average reward and success rate. All of these three metrics robustly demonstrate that Digimon can achieve superior efficacy compared to the SOTA baseline APPLR.
IX-B Ablation Study
@@
Sensitivity. We conduct ablation studies to evaluate the sensitivity of Digimon against various crucial hyper-parameters: (1) high-resistance area sampling threshold $\lambda$ in Algorithm 1 and (2) high-resistance point identification threshold $\eta$ in Algorithm 2.
@@
Failure Trajectory Filtering. We evaluate the necessity of our failure trajectory filtering in Digimon and its advantages for improving navigation performance. The results are shown in Table V-B. We observe that without the failure filtering trajectory mechanism, three out of five metrics deteriorate significantly, particularly the collision rate and success rate, which are critical to the robot’s safety. In contrast, without failure filtering, the NS and ATT metrics show only marginal improvements. This indicates that filtering out failure trajectories can substantially enhance overall navigation performance in terms of safety and task completion rate while not sacrificing total travel time.
@@
High-resistance Area Up-sampling Ratio $\lambda$. We then study the influence of $\lambda$ in Alg. 1 on the navigation performance. Table VIII shows that $\lambda=0.4$ yields the best performance for the navigation system. Varying $\lambda$ introduces fluctuations across various metrics, while all within a 10% range. This demonstrates the necessity of fine-tuning the high-resistance area up-sampling ratio $\lambda$ to balance the proportion of difficult trajectories within the total training set. If $\lambda$ is set too high, the training set will contain too many challenging trajectories, complicating the meta-planner’s ability to understand the complete map and complete the task. If $\lambda$ is set too low, it will revert to the original training method, hindering the meta-planner’s ability to learn from those under-fitted trajectories.
@@
In Table VIII, we show the performance metrics of Digimon over different threshold angles. While there are some improvements in certain metrics (e.g., NS and CR) at higher threshold angles, the performance data clearly converges around 90° as the optimal threshold for overall system performance. This peak suggests at 90°, the system achieves a balance between speed, efficiency, and safety, highlighting a degree of robustness and adaptability. Similar to the high-resistance area up-sampling ratio $\lambda$, the threshold angle should also be carefully tuned to enable the best performance. Thus, we set $\lambda=0.4$ and $\eta=90°$ for all other experiments.
X Potential Social Impacts
@@
The development and deployment of Digimon, aimed at enhancing the performance of meta-planner policy in robotic navigation, promises significant positive social benefits. By enabling robots to navigate more effectively and reliably, our tool can greatly facilitate the integration of robotics into our daily life. For example, in healthcare settings, more efficient navigation can enable robots to deliver medication, assist in surgeries, or provide companionship with greater precision and safety. In homes, robots equipped with advanced navigation capabilities can assist individuals with disabilities, offering them greater independence and quality of life. Furthermore, in disaster response situations, robots that can navigate challenging terrains could save lives by reaching areas that are inaccessible to humans. Overall, by improving the reliability and functionality of meta-planners in robotic navigation systems, Digimon has the potential to make significant contributions to society, improving safety, accessibility, and efficiency in numerous critical areas.

### v3

@@ II Diagnosis Effectiveness
Figure 8: RandSampling v.s. EffiTune. Test on Easy Level.
@@
Figure 9: RandSampling v.s. EffiTune. Test on Medium Level.
@@
Figure 10: RandSampling v.s. EffiTune. Test on Difficult Level.
(a) Average Time
[…]
(b) Average Reward
[…]
(c) Average Success Rate
Figure 11: Average time, reward, and success rate of EffiTune and APPLR for different checkpoints during training.
Table V: Navigation performance under different high-resistance area up-sampling ratios $\lambda$.
@@
Table VI: Navigation performance under different trajectory vector change thresholds $\eta$.
@@
III More Results of Efficiency Comparison
@@
In Figure 11, we present additional metrics for various parameter tuner checkpoints during training, including average traveling time, reward, and success rate. It clearly shows that EffiTune outperforms APPLR across all three metrics. Notably, EffiTune requires fewer training steps to achieve shorter traveling times, highlighting the method’s efficiency. Similarly, within the same amount of training steps, EffiTune can also achieve a higher average reward and success rate. All of these three metrics robustly demonstrate that EffiTune can achieve superior efficacy compared to the SOTA baseline APPLR.
IV Ablation Study
@@
Sensitivity. We conduct ablation studies to evaluate the sensitivity of EffiTune against various crucial hyper-parameters: (1) high-resistance area sampling threshold $\lambda$ in Algorithm 1 and (2) high-resistance point identification threshold $\eta$ in Algorithm 2.
@@
Failure Trajectory Filtering. We evaluate the necessity of our failure trajectory filtering in EffiTune and its advantages for improving navigation performance. The results are shown in Table III. We observe that without the failure filtering trajectory mechanism, three out of five metrics deteriorate significantly, particularly the collision rate and success rate, which are critical to the robot’s safety. In contrast, without failure filtering, the NS and ATT metrics show only marginal improvements. This indicates that filtering out failure trajectories can substantially enhance overall navigation performance in terms of safety and task completion rate while not sacrificing total travel time.
@@
High-resistance Area Up-sampling Ratio $\lambda$. We then study the influence of $\lambda$ in Alg. 1 on the navigation performance. Table II shows that $\lambda=0.4$ yields the best performance for the navigation system. Varying $\lambda$ introduces fluctuations across various metrics, while all within a 10% range. This demonstrates the necessity of fine-tuning the high-resistance area up-sampling ratio $\lambda$ to balance the proportion of difficult trajectories within the total training set. If $\lambda$ is set too high, the training set will contain too many challenging trajectories, complicating the parameter tuner’s ability to understand the complete map and complete the task. If $\lambda$ is set too low, it will revert to the original training method, hindering the parameter tuner’s ability to learn from those under-fitted trajectories.
@@
In Table II, we show the performance metrics of EffiTune over different threshold angles. While there are some improvements in certain metrics (e.g., NS and CR) at higher threshold angles, the performance data clearly converges around 90° as the optimal threshold for overall system performance. This peak suggests at 90°, the system achieves a balance between speed, efficiency, and safety, highlighting a degree of robustness and adaptability. Similar to the high-resistance area up-sampling ratio $\lambda$, the threshold angle should also be carefully tuned to enable the best performance. Thus, we set $\lambda=0.4$ and $\eta=90$° for all other experiments.
V Potential Social Impacts
@@
The development and deployment of EffiTune, aimed at enhancing the performance of parameter tuner policy in robotic navigation, promises significant positive social benefits. By enabling robots to navigate more effectively and reliably, our tool can greatly facilitate the integration of robotics into our daily life. For example, in healthcare settings, more efficient navigation can enable robots to deliver medication, assist in surgeries, or provide companionship with greater precision and safety. In homes, robots equipped with advanced navigation capabilities can assist individuals with disabilities, offering them greater independence and quality of life. Furthermore, in disaster response situations, robots that can navigate challenging terrains could save lives by reaching areas that are inaccessible to humans. Overall, by improving the reliability and functionality of parameter tuners in robotic navigation systems, EffiTune has the potential to make significant contributions to society, improving safety, accessibility, and efficiency in numerous critical areas.